OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = two thirdsx − 10 y = two thirdsx + 10 y = −three halvesx − 10 y = −three halvesx + 10

OpenStudy (anonymous):

\[2x - 3y = 3\] \[2x - 3 = 3y\] \[y = \frac{ 2x - 3 }{ 3 }\] \[m = \frac{ 2 }{ 3 }\] \[m = -\frac{ 3 }{ 2 }\] --------------------- \[y - 2 = -\frac{ 3 }{ 2 } (x + 8)\] \[2y - 4 = -3x - 24\] \[y = \frac{ -3x - 20 }{ 2 }\] \[y = -\frac{ 3 }{ 2 }x - 10\]